A block moves along a horizontal surface with a 0.61 coefficient of kinetic friction. The block is pulled by a rope that exerts a 162 N tension force at 370 above the horizontal. Draw the FBD. Determine the kinetic friction.


1
Expert's answer
2022-04-02T10:10:45-0400

"F_{fr}=F(\sin\alpha-\mu\cos\alpha)=18.6~N."